# Task Aware Dreamer for Task Generalization in Reinforcement Learning

### v4

@@ Appendix D Experimental Details for MuJoCo
We here introduce state-based control tasks, including Half-CheetahFwd-Back, Half-Cheetah-Vel, and Humanoid-Direc-2D, in detail, following the setting of previous meta RL works (Finn, Abbeel, and Levine 2017; Rakelly et al. 2019).
@@
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. For each seed and each task setting, it will take around 1 day. We select 1 as the action repeat for all experiments following.
@@
Table 6: Average cumulative reward (mean $\pm$ one std) over different target region (smaller $\beta$ represents smaller target region and more sparse return) of the best policy trained by Dreamer and TAD in Cheetah__\__Speed. For each $\beta$, we train agents in the train tasks and evaluate them in both train and test environments. Numbers greater than 95 percent of the best performance for each environment are bold.
@@
TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 841.9 $\pm$ 154.9 546.1 $\pm$ 291.6 $7$$\pm$ 170.1 716.8 $\pm$ 136.5
Table 7: Average cumulative reward (mean $\pm$ one std) over different sparse rates of the best policy trained by Dreamer and TAD in Cheetah__\__Speed. For each sparse rate, we train agents in the train tasks and evaluate them in both train and test environments. Numbers greater than 95 percent of the best performance for each environment are bold.
@@ Appendix E Ablation Study
E.1 Ablation Study on Reward Signals
In this part, we do ablation studies on reward signals for model-free methods including CURL and SAC+AE, which both belong to $\Pi_{1}$. We design CURL (w/ r) and SAC+AE (w/ r) by directly adding reward signals into the observation based on CURL and SAC+AE, respectively. As shown in Table. 4, the performance of CURL (w/ r) and SAC+AE (w/ r) is similar to CURL and SAC+AE. The major reason is that CURL (w/ r) and SAC+AE (w/ r)only utilize the observation and reward of the current timestep without historical information. Thus it is still difficult for them to distinguish different tasks.
E.2 Dynamic Generalization
@@
  • Acrobot-Cartpole-Pendulum: includes 7 tasks of artpole-balance, cartpole-balance__\__sparse, cartpole-swingup, cartpole-swingup__\__sparse, acrobot-swingup, acrobot-swingup__\__sparse, and pendulum-swingup. All these tasks aim to control a rod-shaped robot, while they own different embodiments, dynamics, and observations.
@@
  • Cheetah-run-mass($m$): This task distribution is based on the task Cheetah-run and the mass of the robot is $m$ times that of the standard task. Thus different tasks own different dynamics. We train the agents in tasks with $m=0.6,1.0,1.2,1.6$ and test them in tasks with parameters $m=0.8,1.4$.
  • Walker-walk-mass($m$): This task distribution is based on the task Walker-walk and the mass of the robot is $m$ times that of the standard task. Thus different tasks own different dynamics. We train the agents in tasks with $m=0.6,1.0,1.2,1.6$ and test them in tasks with parameters $m=0.8,1.4$.
Then we test Dreamer and TAD in these four settings and report the results. TAD can achieve much greater performance and better convergence compared to Dreamer, as it can better infer the current task. This experiment indicates TAD’s potential in further handling dynamic generalization and even cross-embodiment tasks.
E.3 Sparse Reward
In this part, we will evaluate TAD in more challenging settings with sparse rewards. First, we evaluate Dreamer and TAD in Cheetah__\__speed with different $\beta$, which identifies the region of target speeds. With smaller $\beta$, the reward signals are more sparse since the target intervals are smaller. In the main experiment, we take $\beta=0.2$, and here we evaluate in $\beta=0.2,0.15,0.1$, of which the result is reported in Table 6. For each $\beta$, we take the training parameters ($0.5,\beta$), ($1.5,\beta$), ($2.0,\beta$), ($3.0,\beta$) and test them in tasks with parameters ($1.0,\beta$), ($2.5,\beta$). As shown in Table 6, with the decreasing of $\beta$, the performance of Dreamer and TAD decreases since reward signals are sparse so exploration here is much more difficult. However, our TAD still significantly outperforms Dreamer and shows strong generalization abilities, which shows that TAD can effectively utilize historical information, even sparse rewards.
Moreover, we design Cheetah__\__speed__\__sparse based on Cheetah__\__speed. In Cheetah__\__speed__\__sparse($n$), we make the reward function sparse, i.e., the output reward is the same as Cheetah__\__speed every $n$ timesteps (in step $n-1,2n-1,...$) and 0 otherwise, of which the sparse rate (SR) is $(n-1)/n$. We supplement experiments to evaluate the performance of Dreamer and TAD with $n=5$ (SR=0.8) and $n=10$ (SR=0.9). As shown in Table 7, with the increasing of SR, although the performance of TAD decreases since inferring task context from sparse reward is extremely difficult, TAD still shows strong performance in train tasks and generalizes well to unseen test tasks.
Appendix F Ethics Issues and Broader Impact
Designing agents that can generalize to unseen tasks is a major concern in reinforcement learning. This work focuses on task generalization in reinforcement learning and proposes a novel algorithm Reward Informed Dreamer. One of the potential negative impacts is that algorithms using deep neural networks, which lack interoperability and theoretical guarantee. If we hope to apply them in real-world applications, they may face security and robustness issues, and a possible way is to develop more explainable methods. There are no serious ethical issues as this is basic research. We hope our work can inspire more research on designing agents with stronger generalization abilities.

### v5

@@ Appendix D Experimental details for MuJoCo
We here introduce state-based control tasks, including Half-CheetahFwd-Back, Half-Cheetah-Vel, and Humanoid-Direc-2D, in detail, following the setting of previous meta RL works [27, 29].
@@
Moreover, we introduce some details about our experiments. Our codes are based on Python and the deep learning library PyTorch. All algorithms are trained on one NVIDIA GeForce RTX 2080 Ti. Each seed and each task setting will take around 1 day. We select 1 as the action repeat for all following experiments.
@@
Table 8: Average cumulative reward (mean $\pm$ one std) over different target region (smaller $\beta$ represents smaller target region and more sparse return) of the best policy trained by Dreamer and TAD in Cheetah_\_Speed. For each $\beta$, we train agents in the train tasks and evaluate them in both train and test environments. Numbers greater than 95 percent of the best performance for each environment are bold.
@@
TAD 951.9 $\pm$ 3.3 876.9 $\pm$ 51.1 841.9 $\pm$ 154.9 546.1 $\pm$ 291.6 777.7 $\pm$ 170.1 716.8 $\pm$ 136.5
Table 9: Average cumulative reward (mean $\pm$ one std) over different sparse rates of the best policy trained by Dreamer and TAD in Cheetah_\_Speed. For each sparse rate, we train agents in the train tasks and evaluate them in both train and test environments. Numbers greater than 95 percent of the best performance for each environment are bold.
@@ Appendix E Ablation study
E.1 Dynamic generalization
@@
  • Acrobot-Cartpole-Pendulum: includes 7 tasks of artpole-balance, cartpole-balance_\_sparse, cartpole-swingup, cartpole-swingup_\_sparse, acrobot-swingup, acrobot-swingup_\_sparse, and pendulum-swingup. All these tasks aim to control a rod-shaped robot, while they own different embodiments, dynamics, and observations.
@@
  • Cheetah-run-mass($m$): This task distribution is based on the task Cheetah-run, and the mass of the robot is $m$ times that of the standard task. Thus different tasks own different dynamics. We train the agents in tasks with $m=0.6,1.0,1.2,1.6$ and test them in tasks with parameters $m=0.8,1.4$.
  • Walker-walk-mass($m$): This task distribution is based on the task Walker-walk, and the mass of the robot is $m$ times that of the standard task. Thus different tasks own different dynamics. We train the agents in tasks with $m=0.6,1.0,1.2,1.6$ and test them in tasks with parameters $m=0.8,1.4$.
Then, we test Dreamer and TAD in these four settings and report the results. TAD can achieve much greater performance and better convergence compared to Dreamer, as it can better infer the current task. This experiment indicates TAD’s potential in further handling dynamic generalization and even cross-embodiment tasks.
E.2 Sparse reward
In this part, we will evaluate TAD in more challenging settings with sparse rewards. First, we evaluate Dreamer and TAD in Cheetah_\_speed with different $\beta$, which identifies the region of target speeds. With smaller $\beta$, the reward signals are more sparse since the target intervals are smaller. In the main experiment, we take $\beta=0.2$, and here we evaluate in $\beta=0.2,0.15,0.1$, of which the result is reported in Table 8. For each $\beta$, we take the training parameters ($0.5,\beta$), ($1.5,\beta$), ($2.0,\beta$), ($3.0,\beta$) and test them in tasks with parameters ($1.0,\beta$), ($2.5,\beta$). As shown in Table 8, with the decreasing of $\beta$, the performance of Dreamer and TAD decreases since reward signals are sparse, so exploration here is much more difficult. However, our TAD still significantly outperforms Dreamer and shows strong generalization abilities, which shows that TAD can effectively utilize historical information, even sparse rewards.
Moreover, we design Cheetah_\_speed_\_sparse based on Cheetah_\_speed. In Cheetah_\_speed_\_sparse($n$), we make the reward function sparse, i.e., the output reward is the same as Cheetah_\_speed every $n$ timesteps (in step n1,2n1,n-1,2n-1,...) and 0 otherwise, of which the sparse rate (SR) is (n1)/n(n-1)/n. We supplement experiments to evaluate the performance of Dreamer and TAD with $n=5$ (SR=0.8) and $n=10$ (SR=0.9). As shown in Table 9, with the increasing of SR, although the performance of TAD decreases since inferring task context from sparse reward is extremely difficult, TAD still shows strong performance in train tasks and generalizes well to unseen test tasks.